2. What is the change in temperature of aluminum when a 245 g piece of hot aluminum is dropped into a bin of 87.0 g of water at 2.7 °C and the resulting temperature is 35 °C. (caluminum = 0.900 J/g°C)
1
Expert's answer
2020-06-02T14:08:53-0400

The heat that aluminum transfers to the water heats the water up from its initial to final temperature:


"c_\text{Al}m_\text{Al}\Delta t_\text{Al}=c_\text{w}m_\text{w}(t_\text{w f}-t_\text{w i}),\\
\space\\
\Delta t_\text{Al}=\frac{c_\text{w}m_\text{w}(t_\text{w f}-t_\text{w i})}{c_\text{Al}m_\text{Al}}=\frac{4.2\cdot 87(35-2.7)}{0.9\cdot245}=13.3^\circ\text{C}."